OpenStudy (nuccioreggie):

An oil company fills 1 over 10 of a tank in 1 over 5 hour. At this rate, which expression can be used to determine how long will it take for the tank to fill completely? 1 over 5 ⋅ 10 hours 1 over 10 ⋅ 5 hours 1 over 5 ⋅ 1 over 10 hour 5 ⋅ 10 hours

OpenStudy (nuccioreggie):

I think its B

OpenStudy (michele_laino):

after time t, what is the portion of filled tank?

OpenStudy (nuccioreggie):

B

OpenStudy (nuccioreggie):

Is it C or B

OpenStudy (michele_laino):

we can write this proportion: \[\Large \frac{1}{{10}}:\frac{1}{5} = 1:t\] now use the fundamental property of proportions to find the value of t

OpenStudy (nuccioreggie):

O so i say D

OpenStudy (michele_laino):

what do you get?

OpenStudy (michele_laino):

are you sure?

OpenStudy (nuccioreggie):

Is it D or A lol

OpenStudy (michele_laino):

hint: solving that proportion for t, we get: \[\Large t = \left( {\frac{1}{5} \times 1} \right) \times 10\]

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

In general if we have the subsequent proportion: \[\Large a:b = c:t\] then, applying the fundamental property of proportions, we get: \[\Large t = \frac{{b \times c}}{a}\]

OpenStudy (michele_laino):

in our case, we can write: \[\Large t = \left( {\frac{1}{5} \times 1} \right) \times 10 = \frac{1}{5} \times 10\]

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (nuccioreggie):

@Michele_Laino. A

OpenStudy (michele_laino):

that's right!